# Inter-turbine Modelling of Wind-Farm Power using Multi-task Learning

### v1

@@ 3.3 B-spline Linear Models
Regression spline models are chosen for functions $f_{1}$ and $f_{2}$; these provide greater flexibility in modelling nonlinear outputs compared to polynomial regression [29], whilst still being linear in parameters. Regression splines are piece-wise continuous polynomial functions, joined at “knots”, where the first and second derivatives of adjacent functions are equal; this ensures a smooth transition between polynomials. In practice, polynomials of degree three (cubic splines), balancing smoothness and flexibility, are most commonly used [30] and are chosen here. Specifically, the B-spline basis-function approach is used for its numerical stability and computational efficiency [31]. Figure 2 shows the B-spline basis functions for polynomials of degree three, using four interior knots between zero and one. Parameters (weights) are learned for each basis-function during model fitting, with model predictions calculated as a weighted sum of parameters and basis functions.
Figure 2: B-spline basis functions over the interval of zero and one, with four interior knots.
Using the B-spline basis function approach, the design matrix $\mathbf{X}$ was constructed. The associated decomposed $\mathbf{Q}$ matrix was then used alongside the linear model parameters $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ for the mean ($\boldsymbol{\mu}$) and dispersion ($\boldsymbol{\phi}$) parameters. B-splines of order four were chosen, with four interior knots per feature; this was found to sufficiently capture the shape of the power-curve, with additional knots resulting in no significant improvement in Normalised Mean Squared Error (NMSE) (described in Section 5.3); this is demonstrated in Figure 3 where the NMSE does not significantly improve above four interior knots. Further optimisation of the number of knots for individual features, and the location of knots was not considered, and is left as an area for further model refinement.
Figure 3: Comparison of NMSE score vs. number of interior knots
@@ 3.4 Graphical Model
A graphical representation of the B-spline beta regression (BSBR) model is displayed in Figure 4. Latent variables (to be estimated) are depicted as unshaded circled nodes, whilst the shaded circled node represents the observed variable. The dotted lines indicate parameters linked by a transformation, whilst the arrows indicate that the target parameter is drawn from a distribution using the preceding parameters. The plate represents multiple instances of the contained node.
@@
Here, t relates to the training turbine, whilst $\textit{N}_{t}$ relates to the number of training data points for that turbine. In this initial case, independent models are learned per turbine t, used to model the observed target variable $y_{i,t}$. $\mathbf{y}_{t}$, the column vector with entries $y_{i,t}$ for $i=1,2,\ldots,N_{t}$, is distributed according to the reparameterised Beta distribution,
@@
[…]
@@ 3.5 Inference
=$\mathcal{E}=\left\{\mathbf{y}_{t}\right\}$ (8)
@@
=$\mathcal{H}=\left\{\boldsymbol{\eta}_{t},\boldsymbol{\zeta}_{t}\right\}$ (9)
@@
[…]
@@ 4.1 Hierarchical Bayesian Modelling for Multi-task Learning
In Section 3, the presented BSBR model can be described as being an example of Single Task Learning (STL), since the task-level parameters $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ are learned independently for each task; in this case, there is no-pooling (NP) of information between tasks. In contrast, a model containing parameters solely at the population-level which are shared across all tasks, can be said to have a complete-pooling (CP) modelling structure. In this case, all tasks are treated as identical, which can lead to poor model generalisation to new tasks. Alternatively, hierarchical Bayesian models (also known as multi-level models) provide a middle-ground, containing both task-specific parameters (allowing for differences between population members), and population-level parameters, which encourage task parameters to be similar; this modelling structure can be described as having partial-pooling (PP) of information. These models are often used as an approach to multi-task learning (MTL), where the goal is to improve the performance and generalisation of a model, by training it on multiple related tasks simultaneously [34].
In the context of power-curve modelling, allowing PP of information makes intuitive sense; wind-turbines within a wind-farm are nominally identical, so they are likely to have similar power-curves (at a population-level), but differences may also arise from varying wake patterns at individual turbine locations (the task-level). An additional benefit of PP of information, is that data-poor tasks in particular are able to borrow statistical strength from data-rich tasks. Whilst in this context data availability is not a significant problem, it is commonly in the field of data-driven SHM [16], which has led to these modelling approaches being used as a solution in the domain of Population-based SHM (PBSHM) research [20, 21].
@@ 4.2 Metamodel Design
In the results in Section 6.1, correlations are observed among the learned $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ parameters of the turbine-independent BSBR models, in relation to the turbine $x$ and $y$ spatial coordinates. This fact motivated the integration of a “metamodel”, which can be described as a model over models. This metamodel was designed to capture the inter-turbine relationships in power-curve parameters, which could be used to predict power for previously unobserved turbines, not necessarily in the training data. This model design has a tangible real-world benefit - if one were able to make accurate probabilistic predictions for a variable of interest, at a location without data (and perhaps even without associated telemetry), this may reduce the overall telemetry requirements among a population. Whilst in the use case here data are readily available at each turbine location, it serves as a useful testbed to experiment with the modelling approach.
First-order linear metamodels were chosen here for both simplicity and to reduce the risk of overfitting to the training data. In matrix form, the parameter vectors $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ are described via the linear metamodels,
@@
As in the BSBR NP model, hyperpriors were used to aid model convergence, and were placed over the metamodel parameter matrices assuming normal distributions,
[…]
where $\mathbf{M}_{\boldsymbol{\eta}}=\left\{\mathbf{M}_{\boldsymbol{\eta}_{\mathbf{x}}},\mathbf{M}_{\boldsymbol{\eta}_{\mathbf{y}}}\right\}$ and $\mathbf{M}_{\boldsymbol{\zeta}}=\left\{\mathbf{M}_{\boldsymbol{\zeta}_{\mathbf{x}}},\mathbf{M}_{\boldsymbol{\zeta}_{\mathbf{y}}}\right\}$ and the prior means and standard deviations were we set equal to 0 and 3 respectively, given the distribution of learned parameters seen for the individual BSBR models; these were again found to be weakly informative, while sufficiently aiding convergence. Figure 5 shows a graphical representation of the model. As with Figure 4, arrows represent distribution inputs, whilst dotted lines show calculated inputs, with plates indicating multiple instances of their contained nodes.
@@
Note that the metamodel parameters are located outside the plates, and are therefore learned at the population-level, whilst the parameters $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ are learned specifically for each turbine.
@@ 4.3 Metamodel Predictions
{mdframed}
Following posterior estimation of the population-level metamodel parameters, one can make probabilistic predictions at locations of interest based on spatial coordinates alone,
$\boldsymbol{\eta^{*}}=\mathbf{C_{x}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{x}}}+\mathbf{C_{y}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{y}}}$ (14)
$\boldsymbol{\zeta^{*}}=\mathbf{C_{x}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{x}}}+\mathbf{C_{y}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{y}}}$ (15)
where $\boldsymbol{\eta^{*}}$ and $\boldsymbol{\zeta^{*}}$ are the predicted metamodel parameters for all turbine locations of interest, $\mathbf{{C_{x}}^{*}}$ and $\mathbf{{C_{y}}^{*}}$ are their associated first-order linear regression design matrices for the normalised spatial coordinates, and $\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{x}}}$, $\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{y}}}$, $\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{x}}}$ and $\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{y}}}$ are the estimated metamodel parameters. The predicted parameters can then be used in line with the population-level input features (and associated Q matrix) to make probabilistic power-predictions.
@@ 4.4 Metamodel Benchmarking
The model presented in Section 4 represents a hierarchical Bayesian model (with partial-pooling), incorporating a metamodel—this complete model is simply referred to as the metamodel. To help assess the performance of the metamodel, it is benchmarked against three models which do not account for spatial correlation between turbines. The first of these, is the basic BSBR model shown in Figure 4, which as compared to the metamodel, has no-pooling nor any linear model constraints between turbine-specific parameters. Secondly the metamodel is compared to a similarly partially-pooled hierarchical model (shown in Figure 6), but where the population-level parameters more simply describe the population mean and standard deviation of the turbine-specific parameters, thus not imposing a linear model constraint over them; this should help test the suitability of the linear metamodel assumption over the turbine-specific parameters. For completeness, the metamodel is compared to a complete-pooling approach, where all turbines are considered identical and a single set of model parameters is learned for all turbines; this is shown in Figure 7. For further details, stan code containing the four models is provided in the links in Section 8.
@@
Figure 6: Partially-pooled (PP) BSBR graphical model.
$y_{i,t}$$\boldsymbol{\mu}$$\boldsymbol{\phi}$$\boldsymbol{\eta}$$\boldsymbol{\zeta}$$\boldsymbol{\mu_{{\eta}}}$$\boldsymbol{\sigma_{{\eta}}}$$\boldsymbol{\mu_{{\zeta}}}$$\boldsymbol{\sigma_{{\zeta}}}$$i\in 1:N$
@@ 5.1.1 BSBR Model Demonstration
In the first instance, to demonstrate the suitability of the BSBR model for capturing the heteroscedastic and asymmetric nature of power-curves, the model is trained and tested on individual turbines (i.e. using the no-pooling model definition). A subset of the filtered data that had undergone stratified sampling by yaw angle is used, to provide the models with an approximately even distribution of data across yaw angles, to try to help make the models more generalisable to all wind directions. The training and testing datasets consisted of 4000 and 1000 data points per turbine respectively.
@@ 5.1.2 Metamodel
In the SCADA data, it was noted that there were significant proportions at zero or maximum power; because of this, additional stratified sampling was applied based on measured turbine output power. The reasons for doing this were twofold: (1) to train and assess models equally for all regions of the power-curve (2) because of the heteroscedastic nature of the power-curve, harmonising the proportions of data in different regions of the power-curve between turbines was considered to make comparisons between predictive metrics more reasonable. Since the metamodel aims to capture spatial correlations induced by wake effects, reason (1) is particularly important, since wake effects are more prevalent in the transitional region of the power-curve [23], which is comparatively data-sparse.
Additionally, for the metamodel, since a single model is learned for all data simultaneously, because of increased computational expense a reduced number of training points were needed. The resulting training and testing sets, having undergone stratified sampling on both the yaw angle and output power, consisted of just 250 points per turbine, which were used for the metamodel and benchmark models. Furthermore, since the metamodel and benchmark models are to be trained on a subset of the turbines, as an example 1/4 of the turbines were used, approximately equally spaced throughout the wind-farm. The training and testing turbines are highlighted in Figure 8. The resulting combined training and test sets consisted of 5000 and 20000 data points respectively.
Figure 8: Map showing the training and testing turbines for the metamodel. Red circles show the locations of the turbines used in training, while all turbines (shown in blue) were used in testing.
@@ 5.2 Model Training
As previously discussed, MCMC is used to estimate the posterior distributions of our parameters, given observed levels of power per turbine. MCMC was implemented using stan [27], with 2000 burn-in and 2000 inference samples per chain, and four chains. As discussed in the methods, to aid computational efficiency we use the QR decomposition on the design matrix as recommended in the stan users guide [27]. For the full details of the stan models, please see the code available at: https://github.com/smbrealy/wind_mtl.
@@ 5.3 Model Scoring
In order to quantify the success of each of the models, two performance metrics were used. Firstly, the Normalised Mean Squared Error (NMSE) assesses the goodness-of-fit, or how close point predictions are to their test value. The NMSE functions similarly to percentage error, where a score of zero indicates a perfect model, whilst a score of 100 is equivalent to a model simply predicting the mean value of the data. NMSE is defined as,
@@
Since the models are probabilistic and include predictive variance, the joint-log-likelihood (JLL) is also calculated as a measure of how well the uncertainty in the measured data was captured by the model, which NMSE does not. The JLL is calculated as,
@@
where $f(y_{{i,t}_{s}}|{\mu}_{{i,t}_{s}},{\phi}_{{i,t}_{s}})$ is the probability density function of the beta distribution as defined in equation (1), and S is the number of posterior samples. The larger the JLL, the better the model has captured the true distribution of the test data. Given the prevalence of power-curve modelling in the literature, it’s worth stating that since the model metrics are dependant upon both the training features and test data used, they are best used as a comparison of the models within this paper only.

### v2

@@ 3.3 B-spline Linear Models
Regression spline models are chosen for functions $f_{1}$ and $f_{2}$; these provide greater flexibility in modelling nonlinear outputs compared to polynomial regression [28], while still being linear in parameters. Regression splines are piece-wise continuous polynomial functions, joined at knots, where the first and second derivatives of adjacent functions are equal; this ensures a smooth transition between polynomials. In practice, polynomials of degree three (cubic splines), balancing smoothness and flexibility, are most commonly used [29] and are chosen here. Specifically, the B-spline basis-function approach is used for its numerical stability and computational efficiency [30]. Figure 3 shows the B-spline basis-functions for polynomials of degree three, using four interior knots between zero and one. Parameters (weights) are learned for each basis-function during model fitting, with model predictions calculated as a weighted sum of parameters and basis-functions.
Figure 3: B-spline basis-functions over the interval of zero and one, with four interior knots.
Using the B-spline basis-function approach, the design matrix $\mathbf{X}$ was constructed and used alongside the linear model parameters $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ for the mean ($\boldsymbol{\mu}$) and precision ($\boldsymbol{\phi}$) parameters. B-splines of order four were chosen, with four, uniformly spaced, interior knots per feature. Uniform spacing was chosen for simplicity, while additional knots resulted in no significant improvement in the normalised mean squared error (NMSE) (defined in Section 5.3); this is demonstrated in Figure 4. Further optimisation of the number of knots for individual features, and the location of knots was not considered, and is left as an area for further model refinement.
Figure 4: Comparison of NMSE score vs. the number of (evenly-spaced) interior knots per feature, for the no-pooling BSBR model described in Section 5.1.1.
@@ 3.4 Graphical Model
A graphical representation of the B-spline beta regression (BSBR) model is displayed in Figure 5. Latent variables (to be estimated) are depicted as unshaded circled nodes, while the shaded circled node represents the observed variable. The dotted lines indicate parameters linked by a transformation, while the arrows indicate that the target parameter is drawn from a distribution using the preceding parameters. The plate represents multiple instances of the contained node.
@@
Here, t relates to the training turbine, while $\textit{N}_{t}$ relates to the number of training data points for that turbine. In this initial case, independent models are learned per turbine t, used to model the observed target variable $y_{i,t}$. The column vector $\mathbf{y}_{t}$, with entries $y_{i,t}$ for $i=1,2,\ldots,N_{t}$, is distributed according to the reparameterised beta distribution,
@@
$\{\boldsymbol{\eta_{t}},\boldsymbol{\zeta_{t}}\}\sim\mathcal{N}(0,3)\quad\text{(element-wise)}$ (7)
@@ 3.5 Inference
$\mathcal{E}=\left\{\mathbf{y}_{t}\right\}$ (8)
@@
$\mathcal{H}=\left\{\boldsymbol{\eta}_{t},\boldsymbol{\zeta}_{t}\right\}$ (9)
@@
$=\frac{p(\mathbf{y_{t}}\mid\boldsymbol{\eta}_{t},\boldsymbol{\zeta}_{t})p(\boldsymbol{\eta}_{t})p(\boldsymbol{\zeta}_{t})}{\int\int p(\mathbf{y_{t}}\mid\boldsymbol{\eta}_{t},\boldsymbol{\zeta}_{t})p(\boldsymbol{\eta}_{t})p(\boldsymbol{\zeta}_{t})d\boldsymbol{\eta}_{t}d\boldsymbol{\zeta}_{t}}$
@@ 4.1 Hierarchical Bayesian Modelling for Multi-task Learning
In Section 3, the presented BSBR model can be described as being an example of single task learning (STL), since the task-level parameters $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ are learned independently for each task; in this case, there is no-pooling (NP) of information between tasks. By contrast, a model containing parameters solely at the population-level which are shared across all tasks, can be said to have a complete-pooling (CP) modelling structure. In this case, all tasks are treated as identical, which can lead to poor model generalisation to new tasks. Alternatively, hierarchical Bayesian models (also known as multi-level models) provide a middle-ground, containing both task-specific parameters (allowing for differences between population members), and population-level parameters, which encourage task parameters to be similar; this modelling structure can be described as having partial-pooling (PP) of information. These models are often used as an approach to multi-task learning (MTL), where the goal is to improve the performance and generalisation of a model, by training it on multiple related tasks simultaneously [33].
In the context of power-curve modelling, allowing PP of information makes intuitive sense; wind-turbines within a wind-farm are nominally identical, so they are likely to have similar power-curves (at a population-level), but differences may also arise from varying wake patterns at individual turbine locations (the task-level). An additional benefit of PP of information, is that data-poor tasks in particular are able to borrow statistical strength from data-rich tasks. While in this context data availability is not a significant problem, it is common in the field of data-driven SHM [16], which has led to these modelling approaches being used as a solution in the domain of population-based SHM (PBSHM) research [20, 21].
@@ 4.2 Metamodel Design
In the results in Section 6.1, correlations are observed among the learned $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ parameters of the NP BSBR models, in relation to the turbine $x$ and $y$ spatial coordinates. This fact motivated the integration of a metamodel, which can be described as a model over models. The metamodel was designed to capture the inter-turbine relationships in power-curve parameters, which could be used to predict power for previously unobserved turbines, not necessarily in the training data. This model design has a tangible real-world benefit—if one were able to make accurate probabilistic predictions for a variable of interest, at a location without data (and perhaps even without associated telemetry), this may reduce the overall telemetry requirements among a population. While in the use case here data are readily available at each turbine location, it serves as a useful testbed to experiment with the modelling approach.
First-order linear metamodels were chosen here for both simplicity and to reduce the risk of overfitting to the (potentially noisy) training data. In matrix form, the parameter vectors $\boldsymbol{\eta}$ and $\boldsymbol{\zeta}$ are described via the linear metamodels,
@@
As in the NP BSBR model, hyperpriors were used to aid model convergence, and were placed over the metamodel parameter matrices assuming normal distributions,
$\{\mathbf{M}_{\boldsymbol{\eta}},\mathbf{M}_{\boldsymbol{\zeta}}\}\sim\mathcal{N}(0,3)\quad\text{(element-wise)}$ (13)
where $\mathbf{M}_{\boldsymbol{\eta}}=\left\{\mathbf{M}_{\boldsymbol{\eta}_{\mathbf{x}}},\mathbf{M}_{\boldsymbol{\eta}_{\mathbf{y}}}\right\}$ and $\mathbf{M}_{\boldsymbol{\zeta}}=\left\{\mathbf{M}_{\boldsymbol{\zeta}_{\mathbf{x}}},\mathbf{M}_{\boldsymbol{\zeta}_{\mathbf{y}}}\right\}$ and the prior means and standard deviations were set equal to 0 and 3 respectively, given the distribution of learned parameters seen for the individual NP BSBR models; these were again found to be weakly informative, while sufficiently aiding convergence. Figure 6 shows a graphical representation of the model. As with Figure 5, arrows represent distribution inputs, while dotted lines show calculated inputs, with plates indicating multiple instances of their contained nodes.
@@
Note that the metamodel parameters are located outside the plates, and are therefore learned at the population-level, while the parameters $\boldsymbol{\eta}_{t}$ and $\boldsymbol{\zeta}_{t}$ are determined specifically for each turbine.
@@ 4.3 Metamodel Predictions
Following posterior estimation of the population-level metamodel parameters, one can make probabilistic predictions at locations of interest, based on spatial coordinates alone, $\boldsymbol{\eta^{*}}=\mathbf{C_{x}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{x}}}+\mathbf{C_{y}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{y}}}$ (14) $\boldsymbol{\zeta^{*}}=\mathbf{C_{x}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{x}}}+\mathbf{C_{y}^{*}}\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{y}}}$ (15)
where $\boldsymbol{\eta^{*}}$ and $\boldsymbol{\zeta^{*}}$ are the predicted metamodel parameters for all turbine locations of interest, $\mathbf{{C_{x}}^{*}}$ and $\mathbf{{C_{y}}^{*}}$ are their associated first-order linear regression design matrices for the normalised spatial coordinates, and $\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{x}}}$, $\mathbf{\hat{M}}_{\boldsymbol{\eta}_{\mathbf{y}}}$, $\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{x}}}$ and $\mathbf{\hat{M}}_{\boldsymbol{\zeta}_{\mathbf{y}}}$ are the estimated metamodel parameters. The predicted parameters can then be used in line with the population-level input features, $\mathbf{X}$, to make probabilistic power-predictions.
@@ 4.4 Metamodel Benchmarking
The model presented in Section 4 is a hierarchical Bayesian model, incorporating a metamodel—this model is simply referred to as the metamodel. To help assess the performance of the metamodel, it is benchmarked against three models which have no metamodel structure or spatial component. The first of these, is the NP BSBR model that was first introduced, which treats each turbine independently—i.e. there is no-pooling of information between turbines, with separate sets of parameters for each one. Secondly the metamodel is compared to a partially-pooled (PP) hierarchical model (shown in Figure 7), where the turbine-specific parameters are drawn from shared population-level distributions. For completeness, the metamodel is compared to a complete-pooling (CP) approach, where all turbines are considered identical and a single set of model parameters are learned for all turbines; this is shown in Figure 8. In the PP model, the $\eta$ and $\zeta$ parameter vectors were assumed to be normally distributed, controlled by the population-level distribution parameters $\left\{\boldsymbol{\mu}_{\eta},\boldsymbol{\sigma}_{\eta},\boldsymbol{\mu}_{\zeta},\boldsymbol{\sigma}_{\zeta}\right\}$. For further details, including the details of the hyperpriors placed over the parameters in the models, stan code containing the four models is provided in the links in Section 8.
@@
Figure 7: Partially-pooled ($P$) BSBR graphical model.
$y_{i,t}$$\boldsymbol{\mu}$$\boldsymbol{\phi}$$\boldsymbol{\eta}$$\boldsymbol{\zeta}$ $i\in 1:N$
@@ 5.1.1 BSBR Model Demonstration
In the first instance, to demonstrate the suitability of the BSBR model for capturing the heteroscedastic and asymmetric nature of power-curves, the model is trained and tested on individual turbines (i.e. using the no-pooling model definition). A subset of the filtered data that had undergone stratified sampling by yaw angle is used, to provide the models with an approximately even distribution of data across yaw angles, to try to help make the models more generalisable to all wind-directions. The training and testing datasets consisted of 4000 and 1000 data points per turbine respectively.
@@ 5.1.2 Metamodel
In the SCADA data, it was noted that there were significant proportions at zero or maximum power; because of this, additional stratified sampling was applied based on measured turbine output power. The reasons for doing this were twofold: (1) to train and assess models equally for all regions of the power-curve (2) because of the heteroscedastic nature of the power-curve, harmonising the proportions of data in different regions of the power-curve between turbines was considered to make comparisons between predictive metrics more reasonable. Since the metamodel aims to capture spatial correlations induced by wake-effects, reason (1) is particularly important, since wake-effects are more prevalent in the transitional region of the power-curve [23], which is comparatively data-sparse. The distributions of the resulting (sub)sampled data are shown in Figures 2(a) and 2(b).
Additionally, for the metamodel, since a single model is learned for all data simultaneously, computational expense is increased significantly, so a reduced number of training points were needed. The resulting training and testing sets, having undergone stratified sampling on both the yaw angle and output power, consisted of just 250 points per turbine, which were used for the metamodel and benchmark models. Furthermore, since the metamodel and benchmark models are to be trained on a subset of the turbines, as an example, one quarter of the turbines were used, which were approximately equally-spaced throughout the wind-farm. The chosen training and the testing turbines are highlighted in Figure 9. The resulting combined training and test sets consisted of 5000 and 20000 data points respectively.
Figure 9: Map showing the training and testing turbines for the metamodel. Red circles show the locations of the turbines used in training, while all turbines (shown by the blue markers) were used in testing.
@@ 5.2 Model Training
As previously discussed, MCMC is used to estimate the posterior distributions of our parameters, given observed levels of power per turbine. For all the models in this study, MCMC was implemented using stan [34], with 2000 burn-in and 2000 inference samples per chain, and four chains. In all cases, the maximum $\hat{R}$ convergence statistics did not exceed 1.01, which suggests proper mixing of the chains [35]. Additionally, the minimum effective sample sizes (ESS) were greater than 400, a benchmark considered suitable for stable parameter estimates and proper interpretation of the $\hat{R}$ statistics [35].
@@ 5.3 Model Scoring
In order to quantify the success of each of the models, two performance metrics were used. Firstly, the normalised mean squared error (NMSE) assesses the goodness-of-fit, or how close point predictions are to their test value. The NMSE functions similarly to percentage error, where a score of zero indicates a perfect model, while a score of 100 is equivalent to a model simply predicting the mean value of the data. NMSE is defined as,
@@
Since the models are probabilistic and include predictive variance, the joint-log-likelihood (JLL) is also calculated as a measure of how well the uncertainty in the measured data was captured by the model. The JLL is calculated as,
@@
where $f(y_{{i,t}_{s}}|{\mu}_{{i,t}_{s}},{\phi}_{{i,t}_{s}})$ is the probability density function of the beta distribution as defined in equation (1), and S is the number of posterior samples. The larger the JLL, the better the model has captured the true distribution of the test data. Given the prevalence of power-curve modelling in the literature, it is worth stating that since the model metrics are dependant upon both the training features and test data used, they are best used as a comparison of the models within this paper only.